The quantity supplied of the commodity in the market is 30 units when the

selling price per unit is P7. Derive the supply equation given that because of the increase

in price to P10 per units, the quantity supplied of this commodity increased to 50 units.


1
Expert's answer
2021-11-01T12:05:53-0400

The equation for supply is of the form Q=c+dP.


We first find the slope of the equation which is; "\frac{\delta Q}{\delta P}=\frac{50-30}{10-7}=\frac{20}{3}"


Supply equation then becomes;


Q="c+\frac{20}{3}P"


To find c, plug in any of the price quantity supplied ;


50="c+\frac{20}{3}\times10"


"c=50-66.67=-16.67"


"\therefore" the supply equation becomes;


Q="-16.67" +"\frac{20}{3}P"